OpenStudy (anonymous):

\[\frac{10-20i}{3-i}\]

sam (.sam.):

Multiply the conjugate \[\frac{10-20i}{3-i} \times \frac{3+i}{3+i}\] simplify

OpenStudy (anonymous):

is that 10-20i x 3+i and 3-i x 3+i

sam (.sam.):

yes

OpenStudy (anonymous):

what about the signs being different

OpenStudy (anonymous):

30 -20i^2

OpenStudy (anonymous):

9 + or - i^2

sam (.sam.):

10-20i x 3+i 30+10i-60i-20i^2 30-50i+20 50-50i

sam (.sam.):

i^2=-1

OpenStudy (anonymous):

oh i see, i missed multiplying the 20